Given two dependent random samples with the following results:

Population 1: 44, 39, 42, 31, 40, 36, 42

Population 2: 32, 32,, 27, 18, 38, 30, 37

Use this data to find the 95% confidence interval for the true difference between the population means. Let d=(Population 1 entry)−(Population 2 entry) Assume that both populations are normally distributed.

Find the mean of the paired differences, find the critical value that should be used in constructing the confidence interval, and find the standard deviation of the paired differences, so that you may construct the 95% confidence interval.
Answer #1

Confidence Interval for the paired t test

From the Given Data: Xd = 8.571, Sd = 4.791

The t critical for alpha = 0.05, df = n - 1 = 7-1 = 6 is 2.4469

The Confidence Interval is Given by

Lower Limit = Xd - tcritical * Sd / Sqrt(n) = 8.571 - 2.4461 * 4.791/sqrt(7) = 8.571 - 4.431 = 4.140

Upper Limit = Xd + tcritical * Sd / Sqrt(n) = 8.571 + 2.4461 * 4.791/sqrt(7) = 8.571 + 4.431 = 13.002

The 95% confidence interval is 4.140 13.002

__________________________________________

Calculations Of Mean and SD

Mean = Sum of Observations/n

Standard Deviation = Sqrt(Variance)

Variance = SS/n-1, where SS = SUM (X - Mean)2.

# Difference Mean (X-Mean)2
1 12 8.571 11.758
2 7 8.571 2.468
3 15 8.571 41.332
4 13 8.571 19.616
5 2 8.571 43.178
6 6 8.571 6.61
7 5 8.571 12.752
8 0 0
Total 60 137.714
n 7
Sum 60
Mean 8.571
SS 137.714
Variance 22.9523
Std Dev 4.791